OpenStudy (anonymous):

if a^2 is even....prove that a is even...

OpenStudy (jiteshmeghwal9):

it's known to everyone that the square of a even number is even

OpenStudy (shubhamsrg):

if a = 2k =>a^2 = 4k^2 = 2*something =>even hence proved..

OpenStudy (anonymous):

u r showin proof to..if a is even...show that a^2 is even........:D

OpenStudy (shubhamsrg):

o.O ??

Parth (parthkohli):

It's the same. lol

OpenStudy (jiteshmeghwal9):

\(\LARGE{a^2=even \space number}\) then the number whose square is \(a^2\) is also even

Parth (parthkohli):

If \(a\) is even, then \(a\) is in the form \(2k\). \[2k \times 2k = 4k^2 = 2(2k^2)\]

OpenStudy (shubhamsrg):

maybe your next ques will be: if a is even prove that a+1 is odd!! or even if a is even prove that 2a is even!! ;)

Parth (parthkohli):

Haha, those are axioms! ^

OpenStudy (shubhamsrg):

in "or even" i didnt mean the mathematical meaning of "even" !! :P

OpenStudy (anonymous):

p->q is not the same as q->p.... i.e.,if A=40degrees,..then A is acute..... where as if A is acute then S is 60 degrees.........????:DDDD

OpenStudy (anonymous):

sorry 40 degrees..

OpenStudy (shubhamsrg):

i thing B is 66.75 radians here ??

OpenStudy (shubhamsrg):

nevermind,,am outa here!! :D

Parth (parthkohli):

Didn't I tell the proof?

OpenStudy (jiteshmeghwal9):

haha ! some problem is with this site

OpenStudy (jiteshmeghwal9):

well now look @ this site @anusha.p u will gt ur answer :)

Parth (parthkohli):

Oh yes! I have this proof posted too as a tutorial!

OpenStudy (anonymous):

@jiteshmeghwal9 i want it by direct proof....

OpenStudy (jiteshmeghwal9):

yeah ! these site is providing diffrent types of answers

Parth (parthkohli):

Wait, no. Direct proofing is not possible in this case!

Parth (parthkohli):

Okay, I lied. It actually is... but it's a long one. "To prove: For all whole numbers x, if x2 is even then x is even. A direct proof is difficult, so a proof by contrapositive is preferable." http://en.wikipedia.org/wiki/Proof_by_contrapositive

OpenStudy (jiteshmeghwal9):

& u have gave u the proof using a axiom that the square of a even number is even

Parth (parthkohli):

If you are talking about the \(a = 2k\) proof, then see above.

OpenStudy (jiteshmeghwal9):

Given : \(a^2\) is even then from above axiom the number whose square is \(a^2\) is also even number

Parth (parthkohli):

And then we continue with \(a = 2k\).

OpenStudy (jiteshmeghwal9):

too small proof @anusha.p :D

OpenStudy (jiteshmeghwal9):

well i m also nt sure. am i right @ParthKohli ???

Parth (parthkohli):

Yes.

OpenStudy (jiteshmeghwal9):

k!

OpenStudy (anonymous):

so..wat do u say finally???

Parth (parthkohli):

In this case, if we are given a hypothesis that looks like "If \(a^2\) is even, then \(a\) is even", we may rearrange(only in this case) to make it look like "If \(a\) is even, then \(a^2\) s even too." Presume that \(a = 2k\) where \(k\) is an integer.\[a^2 \implies (2k)^2 \implies 4k^2 = 2(2k^2) \]We know that this is an even number since 2 is a factor.

OpenStudy (anonymous):

can we rearrange......? i mean is it allowed....if so..how??

Parth (parthkohli):

It's just mind-work and logic. Think about it, and don't think about \(p\Rightarrow q\) thing.

OpenStudy (anonymous):

but the actual thing here is the implication only....na

Parth (parthkohli):

"If \(a^2\) is even, then \(a\) is even" is just a fancy way to say "If a is even, then \(a^2\) is even too."

Parth (parthkohli):

Think about it for a while.

OpenStudy (theviper):

\[\Large{a^2=a \times a}\]Since \(a^2\)is even number therefore a is also even because when we multiply anything with an even no. we get an even number ;)

OpenStudy (anonymous):

???

Parth (parthkohli):

@TheViper That's a nice way, but you gotta use these statements as less as possible. "we multiply anything with an even no. we get an even number" is a theorem, and not an axiom. Rest is well :)

OpenStudy (anonymous):

http://uk.answers.yahoo.com/question/index?qid=20101012102515AAlORD8 check this out @parthkolhi

OpenStudy (theviper):

OK thanx @ParthKohli :)

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

@anusha.p Haha, I did give a proof for the contrapositive.

Parth (parthkohli):

Quote, unquote, "I want a direct proof." I told you that we can't just directly prove, and so did Wikipedia. ;P

OpenStudy (anonymous):

so..what should i do.......this quest is given under direct proof secton.....:(

Parth (parthkohli):

Just use my proof. Haha!

Parth (parthkohli):

Okay, don't get the converse. You start with assuming that \(a = 2k\).\[(2k)^2 =4k^2 = 2(2k)^2 \]